### v2

# RobMOT: Robust 3D Multi-Object Tracking by Observational Noise and State Estimation Drift Mitigation on LiDAR PointCloud

Abstract
This work addresses limitations in recent 3D tracking-by-detection methods, focusing on identifying legitimate trajectories and addressing state estimation drift in Kalman filters. Current methods rely heavily on threshold-based filtering of false positive detections using detection scores to prevent ghost trajectories. However, this approach is inadequate for distant and partially occluded objects, where detection scores tend to drop, potentially leading to false positives exceeding the threshold. Additionally, the literature generally treats detections as precise localizations of objects. Our research reveals that noise in detections impacts localization information, causing trajectory drift for occluded objects and hindering recovery. To this end, we propose a novel online track validity mechanism that temporally distinguishes between legitimate and ghost tracks, along with a multi-stage observational gating process for incoming observations. This mechanism significantly improves tracking performance, with a $6.28\%$ in HOTA and a $17.87\%$ increase in MOTA. We also introduce a refinement to the Kalman filter that enhances noise mitigation in trajectory drift, leading to more robust state estimation for occluded objects. Our framework, RobMOT, outperforms state-of-the-art methods, including deep learning approaches, across various detectors, achieving up to a $4\%$ margin in HOTA and $6\%$ in MOTA. RobMOT excels under challenging conditions, such as prolonged occlusions and tracking distant objects, with up to a 59% improvement in processing latency.
Figure 1: A trajectory drift scenario involving a car (ID4) hidden by a van (ID2). (a) At frame 0, four cars are detected and identified (ID1, ID2, ID3, ID4). (b) By frame 34, the method CastTrack [1] incorrectly estimates the position of car ID4 due to state estimation drift. A dashed green rectangle indicates the actual position. (c) Our proposed approach, with its noise mitigation techniques, successfully localizes car ID4 accurately.
3D multi-object tracking (MOT) is one of the pillars of autonomous driving. It is responsible for determining the state of objects in the environment, including their spatial location and motion parameters like velocity and acceleration. Trajectory drift has been observed in state estimation of occluded objects in the top-performing state-of-the-art CasTrack [1], as shown in Figure 1, failing to recover objects under challenging tracking conditions like prolonged occlusion and distant objects as emphasized in Figure 7. Our investigation shows that this drift results from fluctuation noise attached to the incoming 3D detections from deep learning models, which tracking-by-detection methods rely on for object localization on LiDAR point clouds. This noise impacts the precision of state estimation for spatial location and motion parameters of occluded objects that ultimately leads to deviation from the actual motion trajectory.
Another issue is the absence of trajectory legitimacy verification in the MOT literature [1, 2, 3, 4, 5, 6]. These works apply a fine-tuned threshold on detection scores of the detections provided by deep learning models to filter out false positive observations that cause ghost trajectories. This approach is not reliable as it can block observations that belong to legitimate trajectories while leaking false positive observations. Since observations of legitimate trajectories can inherit some of the false positive characteristics, such as low detection scores for distant and partial occlusion scenarios, the MOT methods need to validate trajectories based on the consistency and legitimacy of their historical observations instead.
In this work, we tackle the above-mentioned challenges through Kalman filter refinement to mitigate the trajectory drift noise, Section 2.2, and propose a novel online mechanism to validate and distinguish between legitimate and ghost trajectories based on studying the characteristics of ghost tracks, Section 2.3. Our proposed refinement shows consistency in tracking for occluded objects superior to the latest state-of-the-art approach, CasTrack, demonstrated in Figure 1 and Figure 7. Next, the proposed trajectory validity mechanism incorporates a multi-stage observational gate that allows potential detections of legitimate tracks to pass to the system and an uncertainty formulation used to quantify the temporal improvement of trajectory certainty based on ghost track characteristics. The mechanism shows a $4\%$ and $6\%$ enhancement in HOTA and MOTA metrics, respectively.
We present the above innovative aspects within a 3D MOT tracking framework dubbed RobMOT. Our contribution can be summarized as follows:
  1. 1.
    To the best of our knowledge, this is the first work to uncover the trajectory drift noise associated with detections and its impact on the state estimation of occluded objects in 3D MOT. Our refinement in Kalman filter shows superior performance over the top-performing benchmark, CastTrack [1], on challenging MOT conditions, demonstrated in Figure 7.
  2. 2.
    We propose a novel online trajectory validity mechanism for temporal distinguishing between ghost and legitimate tracks, supported by a multi-stage observational gate for potential observation identification of the legitimate tracks. The mechanism reduces about 80% of ghost tracks and enhances HOTA by 7%, as demonstrated in the ablation study Section 3.2.2.
  3. 3.
    We propose an online 3D MOT framework with exceptional tracking performance, surpassing recent benchmarks to show its robustness across different datasets and detectors. The tracking performance margin of our framework with the state-of-the-art CasTrack [1] reaches $4\%$ in HOTA and $6\%$ in MOTA, and the enhancement in computational latency is up to $53\%$, as demonstrated in Figure 6.
1 Related Work
Figure 2: The RobMOT framework consists of four modules: a multi-stage observational gate for identifying legitimate observations, an association trajectory innovation for trajectories’ association with observations and next-state prediction, a trajectory validity for temporal validation of the trajectories, and a cache for storing and termination.
3D MOT methods on LiDAR point cloud could be categorized into learning and non-learning (classical) approaches. Part of learning-based methods in the literature involves end-to-end deep learning models that combine object detection and tracking [7, 8, 9, 10]. Meanwhile, other methods adapt tracking-by-detection paradigm [11, 12, 13, 14] when the detection of objects is given and the aim is to perform temporal tracking of the objects. The recent research integrates attention mechanism and transformers concepts [15, 16] into MOT to enhance the tracking performance. However, learning-based solutions require evolved hardware to be applicable to real-time applications.
Meanwhile, classical MOT solutions [4, 5, 2, 3, 6, 17, 1] have gained much attention due to their low latency and the comparative performance with learning-based solutions. Despite recent advancements in the classical methods, challenges such as the limitations of the Kalman filter, ghost tracks, and memory management persist, especially in autonomous driving applications where the environment constantly changes. We will shed light on these challenges and discuss the remedies proposed in the literature next.
Kalman filter (KF) is widely used in tracking-by-detection methods [2, 3, 17, 1] for autonomous driving because of its performance in filtering noise from state estimation while dealing with noisy measurements. If the noise is not properly modeled in the filter, it can cause deviation in the state estimation, as the filter’s estimation is sensitive to imprecise measurements. Cao et al. [6] prevent error accumulation in state prediction caused by periodic occlusion of objects by refining the object’s trajectory after object reappearance. However, the object should be recovered in order to refine its trajectory. Moreover, as of the point cloud’s sparsity, the incoming detections from 3D detectors suffer from spatial distortion around the object’s actual location. This spatial distortion results in an imprecise estimation of the motion parameters, ultimately leading to a deviation in the state estimation once the object faces an occlusion.
Ghost tracks are false positive detections generating illusion tracks that could eventually lead to wrong navigation of the AV. Since the tracking-by-detection paradigm relies heavily on detection accuracy, ghost tracks in object tracking can significantly harm the tracking performance. Kim et al. [3] introduce a multi-stage data association method to overcome this issue by employing multiple detection sources. Nevertheless, they still depend on predefined thresholds to prevent false positives. The downside of a fine-tuned threshold while filtering ghost tracks is the potential of leaking false positive detections that exceed the threshold while blocking detections that belong to actual tracks. Other approaches [5, 2] incorporate a filtration step to address false positive detections. Wang.X et al. [2] confirm the legitimacy of a track when three consecutive detections of that track have been observed. In contrast, Wang.L et al.  [5] select objects for the highest 30% detections in the confidence score. However, the assumptions provided by these solutions may not apply to all scenarios, particularly distant objects or objects subject to periodic occlusions.
Memory management plays a critical role in addressing false positive detections. Maintaining ghost tracks in memory can lead to invalid associations. Meanwhile, early elimination may lose legitimate tracks under prolonged occlusion. In the literature, some work [2, 3, 5, 6] set up a threshold to discard tracks from memory based on how long these tracks have not been observed. Although short memory is suitable for ghost track elimination, it could also remove tracks of objects facing prolonged occlusion. On the other hand, long memory will not remove ghost tracks while increasing the computational time. Nagy et al. [17] attempt to address the trade-off by proposing an MOT solution with low latency and leveraging speed to extend the lifetime of objects, thus capturing more occlusion scenarios. While this approach excels in tracking objects with prolonged occlusion, it does not handle false positives. Wu et al. [1] resolve the issue by associating a confidence score to pre-observed objects in object association. This score decreases over time when the object becomes not observed for a period of time to prevent potential associations with false positive detections. Nonetheless, this approach cannot recover objects under prolonged occlusion as their association score will decrease.
This work proposes RobMOT, a framework to mitigate state estimation deviation caused by fluctuation noise. It also provides adaptive identification for ghost tracks with a quick elimination from memory alongside a mechanism that prevents legitimate track detections from being blocked while filtering ghost tracks.
2 Methodology
As shown in Figure 2, the framework consists of four parts. A multi-stage observational gate: responsible for identifying potential observations of legitimate tracks. Refined Kalman filter for noise mitigation: a refined KF that updates the state estimation of cached trajectories and predicts their next state. Trajectory validity: Updates the legitimacy uncertainty of trajectories recently associated to observations. Cache: Holds trajectories stored in the framework and discards terminated tracks.
2.1 Multi-Stage Observational Gate
The multi-stage observational gate is an identification step for potential observations of the legitimate tracks identified by the trajectory validity stage, (see Section 2.3). The primary purpose of this stage is to allow observations of distant and partially occluded objects with low detection scores to enter while preventing overwhelming the system.
There are three types of the observations: observations of existing legitimate tracks, observations of tracks not generated yet by the system, and false observations that cause ghost tracks. Despite ghost tracks identification in the trajectory validity stage, discussed in Section 2.3, a threshold $\alpha_{nconf}$ should be assigned to the incoming observation to prevent overloading the MOT method because 3D deep learning detectors on LiDAR point cloud deliver immense number of detections with various detection scores. Thus, $\alpha_{nconf}$ is assigned a minimal value for detection scores to avoid overloading the system, regardless of the observations’ legitimacy, as this will be assessed later during the trajectory validity stage.
Even though assigning $\alpha_{nconf}$ with a small value allows a high volume of observations to pass to the system, it could block legitimate observations. Hence, We initially identify potential observations of existing legitimate tracks using the same Ecludiean distance threshold $\sigma$ used in the association step (see Section 2.2). Observations located in a distance less than $\sigma$  to a legitimate track, identified by the trajectory validity stage in Section 2.3, are considered legitimate observations. Although the identified potential legitimate observations can pass directly to the system, we have observed that it still causes high latency because of flooding detections from deep learning models on the point cloud. Thus, we provide an additional detection $\alpha_{conf}$ score filtration when $\alpha_{conf}\leq\alpha_{nconf}$ to allow legitimate observations to enter the system while preventing overloading the system. $\alpha_{nconf}$ and $\alpha_{conf}$ are fine-tuned so that no enhancement in the tracking performance is observed since their preliminary purpose is to prevent overloading the system. Lastly, the filtered observations are projected from the LiDAR coordinate to the world coordinate. Algorithm 1 demonstrates these procedures in detail.
Algorithm 1 Multi-Stage Observational Gate
  function MultiStageObservationalGate($D,T,\sigma,\alpha_{c},\alpha_{n}$)
  Input:
   $D$: List of 3D detections from LiDAR
   $T$: List of trajectories in the cache
   $\sigma$: Euclidean distance threshold
   $\alpha_{conf}$: Detection score threshold for potential confirmed observations
   $\alpha_{nconf}$: Detection score threshold for non-confirmed observations
  Output:
  $P$: List of detections that passed the gate
  $P\leftarrow$ []
  for each $det$ in $D$ do
[…]
        continue
     end if
     $confirmed\leftarrow false$
     for each $traj$ in $T$.getConfirmed() do
[…]
        if $dist\leq\sigma$ then
           $confirmed\leftarrow true$
           break
        end if
     end for
[…]
        $detW\leftarrow$ FromLidarToWorld($det$)
        Add $detW$ to $P$
     end if
  end for
  return  $P$
  end function
2.2 Refined Kalman filter for Noise Mitigation
Figure 3: A scenario demonstrates the impact of detection noise on parked cars. Initially, the state estimation of the three cars is positioned correctly in frame 10. The middle frame, frame 17, shows drift in state estimation of cars with ID 0, orange circle, and 2, blue circle, caused by the noise. The state deviation from the original position of the cars is marked by disconnected white line. Considering the proposed noise mitigation (the third scene on the left), the state estimation of cars with ID 2 has no deviation, while the Car 0 deviation has been reduced significantly.
The permitted observations are associated with trajectories cached in the framework. We use the Hungarian algorithm [18] with the algebraic association strategy followed in [17]. The association for an observation is performed by selecting a trajectory with the shortest Euclidean distance between the estimated state and the observation. When the shortest distance to an observation exceeds $\sigma$, a new trajectory is created for the observation.
The framework performs next-state estimation for all trajectories cached in the system. Trajectories associated with an observation will pass through state update and state prediction as demonstrated in Figure 2. Other trajectories will go directly through state prediction. The framework utilizes KF with a constant acceleration motion model.
We have observed drift in state estimation for objects encountering occlusion or leaving the sensor field of view (off-scene). This phenomenon is also observed in the literature, as displayed in Figure 1 when the estimated trajectory of a parked car with ID 4 drifts away from the expected location during the occlusion. Figure 3 shows another scenario when two parked cars with ID 0 and 2 drift from the actual position after being off-scene. Our investigation of the phenomenon concludes that this deviation is caused by trembling in detection bounding boxes around objects’ locations, which results in illusion movement for these objects that eventually impacts state update and prediction in KF.
We model this noise using the algebraic deviation, shown in Equation 1, between the detected bounding boxes $z^{det}_{ij}=(x_{ij},y_{ij})$ and the actual location of the objects $z^{gt}_{ij}=(x_{ij},y_{ij})$ across $k$ observations for $N$ objects using KITTI training dataset.
$\begin{split}\mu_{x}&=\frac{1}{\sum_{i=1}^{N}k_{i}}\sum_{i=1}^{N}\sum_{j=1}^{k_{i}}\left(z^{gt}_{ij}(x)-z^{det}_{ij}(x)\right)\\ \sigma_{x}^{2}&=\frac{1}{\sum_{i=1}^{N}k_{i}}\sum_{i=1}^{N}\sum_{j=1}^{k_{i}}\left(\left(z^{gt}_{ij}(x)-z^{det}_{ij}(x)\right)-\mu_{x}\right)^{2}\end{split}$ (1)
Figure 4: The distribution of the fluctuation noise associated with VirConv [19] detections. The distribution shows Gaussian noise across five different detectors [19, 20, 21, 22, 23].
As depicted in Figure 4, the error distribution shows a Gaussian shape for the noise of the VirConv detector across $x$ and $y$ LiDAR coordinates, which facilitates its integration into KF. Accordingly, we derive the $2\times 2$ covariance matrix $D_{t}$ representing the noise impact on the detections (measurements). Notably, the noise varies from one detector to another as to the variance in detection performance; hence, $D_{t}$ should be recomputed as the employed detector changes. By applying Equation 1 on $y$ coordinate, we obtain a covariance matrix of the noise, Equation 2.
$\mathbf{D_{t}}=\left[\begin{matrix}\sigma_{x}^{2}&0\\ 0&\sigma_{y}^{2}\end{matrix}\right]$ (2)
Since this noise is attached to the incoming detections $z_{t}$ at teim $t$, it harms the computation of the residual term $\hat{y}_{t}$, which is the difference between the current measurement $z_{t}$ and state estimation $\hat{x}_{t|t-1}$ predicted at $t-1$. Consequently, the noise impact on the residual term leads to incorrect updates for the motion parameters of the object. The innovation covariance $S_{t}$ is responsible for weighting the innovation residual $\hat{y}_{t}$ through the Kalman gain $K_{t}$. Therefore, the covariance matrix $D_{t}$ needs to be added to the innovation covariance $S_{t}$ to compensate for the impact of noise on the updated state. Hence, the updated state estimation $\hat{x}_{t|t}$ will consider the impact of the deviation noise on the observation.
Innovation Residual: $\tilde{y}_{t}=z_{t}-H_{t}\hat{x}_{t|t-1}$ (3)
Innovation Covariance: $S_{t}=H_{t}P_{t|t-1}H_{t}^{T}+R_{t}+D_{t}$
Kalman Gain: $K_{t}=P_{t|t-1}H_{t}S_{t}^{-1}$
State Estimation Update: $\hat{x}_{t|t}=\hat{x}_{t|t-1}+K_{t}\tilde{y}_{t}$
Estimation Covariance Update: $P_{t|t}=(I-K_{t}H_{t})P_{t|t-1}$
The main reason for not using $R_{t}$ instead is because it models the incorrectness of the measurement for the point cloud from LiDAR with respect to the real world, while $D_{t}$ models the incorrectness of the detection process with respect to the LiDAR point cloud, which means they are independent. Equation 3 demonstrates the updated state of KF after adapting $D_{t}$ noise.
$H_{t}$ is the observation model to map from estimation to measurement space at time $t$, while $P_{t|t-1}$ is the estimation uncertainty at time $t$ given measurement at $t-1$. In Section 2.4, we will describe how $P_{t|t}$ contribute to trajectory termination in the cache.
2.3 Trajectory Validity
The objective is to monitor the validity (legitimacy) of the existing trajectories in the system while identifying invalid ones, called ghost tracks. Each track has a certainty score that represents its legitimacy. This score is updated whenever a new observation of the track is available. The certainty score changes based on the likelihood of the trajectory being a ghost track. By observing the characteristics of ghost tracks, the ghost tracks are usually formed from intermittent sequences of observations with low to moderate detection scores. Based on these characteristics, the certainty score for track $i$ changes when a new observation is associated with the track at time $t$. The certainty score changes according to the detection score $s_{i}^{t}$ and the duration of absence $d^{i}_{t}$ since the last observation associated with the track.
[…]
$i\text{ at time }t$
$d^{i}_{t}:\text{absence duration since the last associated observation at time }k$
[…]
$f^{i}_{t}(s^{i}_{t},d^{i}_{t}):\text{certainty score of trajectory }i\text{ at time }t$
$f^{i}_{k}(s^{i}_{k},d^{i}_{k}):\text{certainty score of trajectory }i\text{ at time }k$
$f^{i}_{t}(s^{i}_{t},d^{i}_{t})$ $=s^{i}_{t}e^{-d^{i}_{t}}-\frac{d^{i}_{t}}{s^{i}_{t}}+f^{i}_{k}(s^{i}_{k},d^{i}_{k}),\text{ where }s^{i}_{t}>0$ (4)
$d^{i}_{t}$ $=t-(k+1)$
As shown in Equation 4, the formulation consists of three terms: Confidence decay term $s^{i}_{t}e^{-d^{i}_{t}}$ (reward), Absence duration term (penalty) $\frac{d_{t}}{s^{i}_{t}}$, and Temporal validity $f^{i}_{k}(s^{i}_{k},d^{k}_{t})$ term.
2.3.1 Confidence decay
The term is responsible for increasing the certainty score of track $i$ as the detection score $s^{i}_{t}$ of the associated observation increases as a reward, indicating the detection model’s confidence in this observation. However, ghost tracks with moderate detection scores can receive high rewards, potentially misleading the validation mechanism. Thus, a decay rate,$e^{-d^{i}_{t}}$, is attached to the detection score $s^{i}_{t}$. The decay $e^{-d^{i}_{t}}$ increases as the duration $d^{i}_{t}$ of the current observation at time $t$ to the previous observation at time $k$ increases. Given the intermittent characteristic of ghost tracks, the decay term decreases the detection score for potential ghost tracks in the system.
2.3.2 Absence duration
The penalty term $\frac{d^{i}_{t}}{s^{i}_{t}}$ lowers the certainty score as the absence duration $d^{i}_{t}$ of a new observation for the track increases. This term will diminish the impact of the reward with an extended absence duration that indicates a high potential of being a ghost trajectory. However, this term can harm legitimate tracks that encounter occlusion. Since observations of legitimate tracks tend to have high detection scores $s^{i}_{t}$, we include inverse proportional to the detection score to decrease the impact of this term on the legitimate tracks.
2.3.3 Temporal validity
The inclusion of $f^{i}_{k}(s^{i}_{k},d^{i}_{k})$ integrates the previous certainty score calculated at time $k$ to the current certainty score $f^{i}_{t}(s^{i}_{t},d^{i}_{t})$ of time $t$. This term adds temporal validation for trajectories in the system by incorporating historical certainty into the current certainty score so that $f^{i}_{t}(s^{i}_{t},d^{i}_{t})$ represents a certainty score of all trajectory states.
Figure 5: A graphical visualization of the certainty scoring formula over detection scores of range $[0,5]$ and absence duration $[0,5]$ frames. The red point is the highest certainty score when the detection score is highest while the absence duration is lowest. In contrast, the blue point is the lowest certainty score when the detection score is at the lowest value, while the absence duration is the highest. The surface shows the remaining values.
A visual representation of the distribution of Equation 4 is highlighted in Figure 5; the relation between the detection score, the absence duration, and its impact on the certainty score.
This mechanism monitors the validity of trajectories in the system individually until a certainty score of a trajectory exceeds a pre-defined threshold $\alpha_{legit}$ that indicates the trajectory is legitimate. In this case, the validation mechanism changes the trajectory status from “not confirmed” to “confirmed” and terminates the validation mechanism to this trajectory. Thus, confirmed trajectories cannot be changed to “not confirmed" as the mechanism is terminated for these tracks. Algorithm 2 demonstrates trajectory validation procedures in detail.
Algorithm 2 Trajectory Validity
  function UpdateTrajectoryCertainty($T,\alpha_{legit}$)
  Input:
   $T$: List of trajectories in the cache
   $\alpha_{legit}$: Trajectory legitimacy threshold
  for each $traj$ in $T$ do
     if $traj$.hasNewObservation() then
        $s_{t}\leftarrow traj$.getObservation().detectionScore()
        $d_{t}\leftarrow traj$.getAbsenceDuration()
        $f_{k}\leftarrow traj$.getCertaintyScore()
        $f_{t}=s_{t}e^{-d_{t}}-\frac{d_{t}}{s_{t}}+f_{k}$
        $traj$.setCertaintyScore($f_{t}$)
        if $traj$.getCertaintyScore() $\leq\alpha_{legit}$ then
           $traj$.setValidityStatus(true)
        end if
     end if
  end for
  end function
2.4 Trajectory Termination and Caching
All trajectories are saved in a cache module. In the literature, a trajectory is terminated or discarded when there are no further observations for that trajectory for several frames. Generally, it is not trivial to define an absence threshold as it could risk eliminating objects under prolonged occlusion or maintaining ghost tracks for a long. Instead, in our framework, we discard the trajectory when the estimation uncertainty, $P_{t|t}$ in Equation 3, becomes high, as it can lead to mismatching in object association. We have observed that ghost tracks tend to have limited observations, leading to high uncertainty in state estimation that matches the intended goal. Meanwhile, legitimate tracks maintain low estimation uncertainty, even those under occlusion. Thus, estimation uncertainty is used to terminate tracks from the cache. Algorithm 3 demonstrates the termination procedures in the cache.
Algorithm 3 Trajectory Termination
  function trajectory Termination($T,\alpha_{cov}$)
  Input:
   $T$: List of trajectories in the cache
   $\alpha_{cov}$: Estimation uncertainty threshold to terminate a trajectory
  for each $traj$ in $T$ do
     $estCertaintyX\leftarrow traj.$getCovariance_P(0)
     $estCertaintyY\leftarrow traj.$getCovariance_P(1)
     $discard\leftarrow$ false
     if $estCertaintyX>\alpha_{conv}$ then
        $discard\leftarrow$ true
     end if
     if $estCertaintyY>\alpha_{conv}$ then
        $discard\leftarrow$ true
     end if
     if $discard$ then
        Discard $traj$ from $T$
     end if
  end for
  end function
3 Results and Discussion
The framework is evaluated using two benchmark datasets for autonomous cars: KITTI [24] and Waymo Open Dataset [25] (WOD). The KITTI dataset consists of $50$ sequences, split into $21$ training sequences and $29$ testing sequences recorded around Karlsruhe, Germany, that involve a variety of challenging conditions. Meanwhile, WOD contains about $1150$ segments recorded in various locations across the United States. The experiments are conducted on a machine with a processor AMD Ryzen 9; no GPU is involved since the framework only utilizes the CPU.
The evaluation of the proposed approach consists of an extensive quantitative and qualitative evaluation with recent benchmarks in this section and a sensitive analysis of each proposed component to evaluate its contribution to the tracking performance, which will be covered in the ablation study, Section 3.2.
3.1 Evaluation with Benchmarks
This section consists of quantitative tracking performance evaluation with the latest state-of-the-art MOT methods on KITTI and WOD datasets and qualitative evaluation focusing on challenging occlusion conditions with the top-performing benchmark selected from the quantitative evaluation.
Figure 6: Tracking performance comparison on KITTI dataset validation split with CasTrack [1] across five different detectors: VirConv [19], CasA [20], PointRCNN [22], PV-RCNN [23], and Second [21]. RobMOT (in red) consistently outperforms CasTrack (in blue) across all detectors, with significant improvements in both HOTA and MOTA metrics, particularly with the PV-RCNN detector, and a substantial reduction in IDSW. Additionally, RobMOT demonstrates a minimum 40% increase in processing speed (fps) across all detectors, highlighting its efficiency and robustness. The percentage improvements are indicated in green.
3.1.1 Quantitative Evaluation
Table 1: Tracking performance on KITTI dataset testing split with the latest MOT benchmarks
Method HOTA MOTA AssA AssRe IDSW CasTrack* [1] 81% 91.9% 84.2% 87.6% 24 PC-TCNN [26] 80.9% 91.7% 84.1% 87.5% 37 Rethink MOT [5] 80.1% 91.5% 83.6% 87.6% 46 MSA-MOT [27] 78.5% 88.0% 82.6% 85.2% 91 DeepFusion-MOT [2] 75.5% 84.6% 80.1% 82.6% 84 PC3T [28] 77.8% 88.8% 81.6% 84.8% 225 Mono-3D-KF [29] 75.5% 88.5% 77.6% 80.2% 162 PolarMOT [30] 75.2% 85.1% 76.95% 80.0% 462 TripletTrack [31] 73.58% 84.32% 74.66% 77.3% 322 RobMOT* (Ours) 81.22% 90.5% 85.8% 89.7% 7
    Methods marked with * use the same detector
Initially, we evaluate the tracking performance on the KITTI testing dataset with the latest state-of-the-art tracking methods listed in the KITTI leaderboard. As illustrated in Table 1, our framework shows a dramatic drop in the ID switch metric (IDSW) compared to the benchmarks, with a difference of 17 IDs, which shows the tracking stability of the proposed approach. Additionally, the framework outperforms in the Higher Order Tracking Accuracy (HOTA) and association accuracy (AssA) metrics with margins of $0.22\%$ and $1.6\%$, respectively. However, CasTrack is superior in Multi-object Tracking Accuracy (MOTA) metrics with a margin of $1.4\%$. Although HOTA measures the consistency of the tracking performance by combining association, localization, and detection compared to MOTA [32], Table 1 shows comparable performance between the proposed solution and CasTrack [1] in HOTA and MOTA metrics despite the frequent change in ID during tracking for CasTrack that reaches $24$ times.
Another evaluation is conducted on WOD testing dataset. The dataset consists of simple to moderate tracking conditions labeled “Level 1" and challenging conditions labeled “Level 2" in Table 4. Our proposed solution shows superior tracking performance on the recent benchmark, including CasTrack [1], with a reduction in the false positives ratio (FP).
Table 2: Tracking performance on Waymo Open dataset testing split with the latest MOT benchmarks
Level 1 Level 2
[…]
CasTrack* [1] 66.95% 7.74% 63.66% 7.36%
ImmotralTrack [33] 63.77% 8.94% 60.55% 8.48%
SimpleTrack [34] 63.53% 8.66% 60.30% 8.75%
CenterPoint [35] 62.58% 9.03% 59.38% 9.41%
AlphaTrack [36] 58.86% 10.10% 55.67% 9.55%
RobMOT* (Ours) 67.02% 7.60% 63.67% 7.22%
    Methods marked with * use the same detector
Despite the superior performance that our approach achieves in KITTI (Table 1) and WOD (Table 2) datasets, CasTrack [1] shows close tracking performance. Thus, we perform a sensitive and qualitative analysis with CasTrack [1] to show the capability of the two methods under different detectors and challenging tracking conditions.
The performance of tracking-by-detection approaches is highly dependent on the utilized detector. High detection performance enhances tracking performance, making evaluating a tracking approach a tricky task. To ensure the superiority of a tracking method, it should maintain its outperformance across various detectors that confirm the contribution of the tracking method. To this end, we conduct a sensitivity analysis across five detectors with different detection performances on KITTI evaluation dataset and compare the proposed approach with CasTrack [1]. Figure 6 shows the tracking performance across the detectors using three different metrics: MOTA, HOTA, and IDSW. Although both methods show close performance in HOTA and MOTA with VirConv [19], CasA [20], and PointRCNN [22], CasTrack [1] performance drops dramatically in PV-RCNN [23] with a margin of more than $6\%$ in MOTA and $4\%$ in HOTA. Meanwhile, our proposed method shows superiority across almost all detectors while maintaining the lowest IDSW compared to CasTrack [1]. These results validate the contribution of the proposed approach while maintaining superior performance across different detectors. In addition, the proposed approach achieved higher frame per second (fps) than CasTrack [1] across the detectors, with enhancement in fps reaching $+59\%$ with the Second [21] detector.
Figure 7: A qualitative evaluation of our framework performance with CasTrack [1] in two challenging multi-object tracking scenarios by projecting 3D tracking results into 2D frames. The first two rows show a scenario where a distant car (ID:41) faces multiple occlusions and distortion in detection. The second two rows show a scenario of a car (ID:3) occluded by a building for about 50 consecutive frames. Our framework tracks the distant car in the first scenario under distortion in the observation and occlusions while providing precise state estimation of the prolonged occluded car in the second scenario with successful recovery after the occlusion.
3.1.2 Qualitative Evaluation
Although the quantitative evaluation shows the overall performance of a tracking approach, it is not sufficient to evaluate the method under challenging tracking conditions such as occlusions. Thus, we select two challenging conditions to emphasize the contribution of the proposed approach. The first condition represented in the first two rows in Figure 7 is for a distant car that faces frequent occlusions for about 90 frames. The detector is unified for both methods, ours and CasTrack [1], to measure the contribution of the tracking approach. Our proposed solution maintains tracking for the car with ID $41$ to frame $193$ when the car leaves the field of view of LiDAR, while CasTrack [1] lost the car at frame $180$.
Another challenging scenario is illustrated in Figure 7 in the third and fourth rows. In frame $57$, a distant car behind a tree is occluded by a building in the following frames until frame $110$, when the car partially appears again. Our proposed method shows robustness in state estimation of the car during occlusion, as shown in frames $102$ and $107$, and successfully reassociates the car when it appears in frame $110$. On the other hand, CasTrack [1] loses track of the car during the occlusion and considers the re-appearance of the car as a new car associated with a new trajectory.
The two scenarios show the robustness of the proposed approach and capability to handle challenging occlusion scenarios compared to the recent state-of-the-art CasTrack [1].
3.2 Ablation Study
This section will cover a sensitivity analysis conducted on the proposed method to show the contribution of each component proposed in this work and quantitatively measure its impact on tracking performance.
3.2.1 Trajectory Drift
This research raises concerns about noise injected in the incoming detections, which leads to deviation in the state estimation of objects during occlusion. The noise is observed in multiple scenarios (Figure 3) and the recent state-of-the-art method CasTrack [1] (Figure 1), which emphasizes the existence of this problem. This section demonstrates the quantitative impact of this noise on the tracking performance and the contribution of the proposed noise mitigation approach discussed in this work in improving the performance by mitigating the noise.
Table 3: Performance comparison of different detectors with and without the proposed refinement for noise mitigation $D_{t}$.
Metric $D_{t}$ VirConv CasA PointRCNN PV-RCNN Second HOTA Y 83% 80.5% 75.7% 78.1% 76.6% N 80.5% 77.2% 72.1% 73.4% 73.5% Improvement - +2.5% +3.3% +3.6% +4.7% +3.1% MOTA Y 87.2% 86.4% 81.5% 84.2% 83.2% N 82.4% 80.8% 76% 77.6% 78.8% Improvement - +4.8% +5.6% +5.5% +6.6% +4.4% IDSW Y 8 4 11 9 8 N 11 19 40 44 33 Improvement - -3 -15 -29 -35 -25 MT Y 479 509 501 496 492 N 470 498 477 484 483 Improvement - +9 +11 +24 +12 +9
In this experiment, we compare the performance of the Kalman filter without the proposed refinement for noise mitigation, $D_{t}$, and the contribution by including the refined Kalman filter. The experiment is conducted on five different detectors to justify the contribution of the proposed refinement Kalman filter, as demonstrated in Table 3. The comparison focuses on metrics MOTA, HOTA, IDSW, and the number of mostly tracked objects (MT). Since our method is not a learning approach, the evaluation is conducted on the KITTI training and evaluation dataset of $21$ streams in total. The proposed refinement in Kalman filter consistently enhances the tracking performance across the five detectors, as shown in Table 3. The improvement in HOTA reached $4.7\%$ with PV-RCNN [23] detector and $6.6\%$ in MOTA. Moreover, it shows consistent state estimation for trajectories that lead to fewer ID switches (IDSW), reduction by $35$ IDs, and more tracked objects (MT); about $24$ objects increased with the PointRCNN [22] detector.
These results, including the qualitative results represented in this work, confirm the contribution of the proposed Kalman filter refinement that mitigates the noise associated with the detections, enhancing state estimation and tracking performance.
3.2.2 Track Validity and Ghost Tracks
The second concern raised in this research is ghost tracks. Ghost tracks are a form of consecutive false positive observations that generate illusion trajectories that could mislead MOT methods. This work proposes an online trajectory validation mechanism to identify ghost tracks among legitimate trajectories in the system. We validate the contribution of the proposed mechanism in two steps. First, we demonstrate a qualitative scenario where ghost and legitimate tracks exist and the capability of the mechanism to identify legitimate tracks. Second, we perform a sensitivity analysis by disabling the mechanism and quantifying the improvement in the tracking performance.
Figure 8 reports a qualitative experiment of the trajectory validation mechanism for identifying legitimate tracks. The figure shows a scenario of three legitimate tracks for three cars with ID $49$, $48$, and 00 and two ghost tracks with ID $50$ and $48$ at frame $277$. A car with ID $47$ is identified as a legitimate track at frame $279$, while the car with ID $49$ is confirmed at frame $285$. The ghost tracks with ID $49$, $48$, and 00 remain unconfirmed until discarded from the system. This scenario shows the capability of the mechanism to identify legitimate tracks among ghost tracks.
Figure 8: This scenario illustrates the proposed trajectory validity mechanism’s capability of distinguishing between real tracks (Car ID:47, Car ID:49) and ghost tracks (Car ID:48, Car ID: 50). It showcases the ability to accurately identify and maintain a ’Not Confirmed’ status for the ghost track while recognizing Car ID:49 and Car ID:47 as legitimate tracks.
Table 4: Performance evaluation of our framework with and without the proposed trajectory validity mechanism across five detectors
Detector Traj. Validity HOTA $\uparrow$ MOTA $\uparrow$ IDSW $\downarrow$ PointRCNN [22] N 69.38% 63.59% 58 PointRCNN [22] Y 75.66% 81.46% 11 Improvement - +6.28% +17.87% -47 PV-RCNN [23] N 73.92% 73.72% 65 PV-RCNN [23] Y 78.13% 84.2% 9 Improvement - +4.21% +10.45% -56 Second [21] N 73.69% 75.78% 39 Second [21] Y 76.64% 83.22% 8 Improvement - +2.95% +7.44% -31 CasA [20] N 78.28% 80.75% 26 CasA [20] Y 80.53% 86.36% 4 Improvement - +2.25% +5.61% -22 VirConv [19] N 80.66% 83.03% 21 VirConv [19] Y 83.01% 87.17% 8 Improvement - +2.35% +4.14% -13
To support the proposed trajectory validity mechanism’s impact on tracking performance, we disable the mechanism and quantitatively evaluate the tracking performance on the training and evaluation KITTI dataset. To demonstrate the generalization of the proposed mechanism, we employ five different detectors and quantify the enhancement in performance across five detectors when the trajectory validity mechanism is enabled. While the trajectory validity mechanism is disabled, we employ threshold on the detection score as proposed and fine-tuned in CasTrack [1] work. The evaluation involves HOTA, MOTA, and IDSW metrics to represent the overall tracking performance. Table 4 shows that enabling the trajectory validity mechanism improves HOTA across all detectors, with a margin reaching $6.28\%$ with PointRCNN [22] detector. Moreover, there is a significant improvement in the multi-object tracking accuracy (MOTA) metric for overall detectors that peak with the PointRCNN detector with up to $17.87\%$ enhancement in MOTA. Since the trajectory validity mechanism identifies legitimate tracks while allowing their observations to pass through the system, the tracking performance becomes more consistent with less misidentification for objects that reflect few IDSW values, as shown in Table 4. It results in a reduction in IDSW reaches $56$ with PV-RCNN [23] detector.
The combination of the qualitative results in Figure 8 and the quantitative results in Table 4 confirms the contribution of the proposed mechanism to object tracking performance.
4 Conclusion
This paper discusses two critical challenges in multi-object tracking (MOT): state estimation deviation and ghost tracks. Our research reveals that the noise associated with detections from the detector affects the accuracy of bounding box localization, leading to errors in estimating the motion parameters of these objects. This problem causes trajectory estimation to drift, particularly when objects encounter occlusion, and has been noted in recent advanced MOT methods. Another type of noise is formed from false positive observations that lead to ghost tracks. While the literature relies on fine-tuned thresholds on the detection score to filter observations contributing to these tracks, it still leaks false observations and blocks observation of legitimate tracks, eventually impacting tracking performance. We addressed this issue with a refined Kalman filter compensating for trajectory drift noise, allowing drift mitigation, and recovering prolonged occluded objects. The work also introduces a novel online trajectory validity mechanism to identify legitimate tracks based on ghost track characteristics in MOT methods.
The proposed solutions are combined in an online MOT framework, RobMOT, that outperforms state-of-the-art trackers in KITTI and Waymo Open Dataset. The framework shows superior tracking performance across various detectors with a margin reaching $4\%$ in HOTA and $6\%$ in MOTA with the top-performing MOT benchmark. The framework qualitatively shows its substantial tracking performance for distant and recovering prolonged occluded objects when SOTA fails. An ablation study has been conducted to demonstrate the improvement in the tracking performance resulting from each component proposed in this work. The study shows that mitigation of the state estimation drift noise enhances HOTA by $4.7\%$ and MOTA by 6.6%. Meanwhile, the proposed trajectory validity mechanism significantly improves the tracking performance, reaching $6.28\%$ in HOTA and 17.87% in MOTA, dramatically reducing the IDSW metric to $56$ IDs. The framework shows its applicability to real-time solutions with CPU constraints and high computational processing.
Appendix A Thresholds and Parameter Values
Table 5: List of Thresholds and Parameter Values
Detector $D_{t}(\sigma_{x}^{2})$ $D_{t}(\sigma_{y}^{2})$ $\sigma_{nconf}$ $\sigma_{conf}$ $\sigma_{legit}$ $\sigma_{conv}$ VirConv [19] 0.016629 0.005334 0 -1 20 4 CasA [20] 0.030696 0.015416 0 0 25 4 PointRCNN [22] 0.032043 0.009945 0 0 35 4 PV-RCNN [23] 0.034076 0.012463 0.5 0.5 20 4 Second [21] 0.037623 0.013561 -1 -2 10 4